# The Completeness of Eigenstates in Quantum Mechanics

### v1

@@ Abstract
The Schrödinger equation (SE) is the fundamental dynamical equation describing quantum systems, and its linearity directly gives rise to the superposition principle. The completeness of a set of orthonormal eigenstates precisely characterizes this superposition principle. Completeness requires that any state of the system can be expressed as a linear combination of these eigenstates; conversely, any linear combination of the eigenstates also constitutes a possible state of the system [Dirac]. This property has been intuitively verified in numerous Young’s double-slit experiments [Young] and enables the prediction of quantum jump probabilities [jump].Born’s probability interpretation [Born] further endows eigenstates with physical significance: the squared modulus of the coefficient of each eigenstate in a superposition corresponds to the probability of the particle being in that state. The physical meaning of completeness is interpreted as a representation transformation [Transformation].
@@
The time-independent SE is a type of (generalized) Sturm-Liouville (S-L) equation. Mathematical research on the completeness of solutions to S-L equations dates back to the Hilbert-Schmidt theory in 1907 [Hilbert]. In 1910, Weyl extended S-L equations from finite to infinite intervals. Courant and Hilbert proved the completeness of the discrete spectrum (bound states) in 1924 [discrete]. Titchmarsh systematically developed the eigenfunction expansion theory for singular S-L operators in 1946, establishing the completeness of the continuous spectrum (free states) and mixed spectrum (coexisting bound and free states) under certain conditions [Titchmarsh]. Kodaira comprehensively summarized previous results in 1949 and developed simplified proof methods for completeness [Kodaira][Well]. Since then, progress in completeness research has been slow—no general proof for the completeness of S-L equations has been achieved to date. Additionally, mathematical proofs of completeness are complex and lack clear physical interpretation.
@@
For quantum mechanics, however, it suffices to prove the completeness of SE solutions. Furthermore, the potential function can be restricted to forms encountered in physical scenarios, eliminating the need for generalized discussions. We first classify potential functions into two categories: symmetric and asymmetric potentials (Sec. III and IV). Following Titchmarsh’s classification method [Titchmarsh1], we further divide these two categories into 8 types based on four possible asymptotic behaviors of the potential at infinity (infinite, constant, finite oscillatory, and infinite oscillatory). Among these 8 types, rigorous theoretical proofs exist for only a small number of cases (Sec. II.1). For each unproven case, we provide proofs through multiple representative examples—some via rigorous theoretical deductions and others via numerical simulations. In these proofs, the spectral function retains physical meaning (momentum spectrum), and the form of completeness aligns with the interpretation of representation transformation (coordinate and momentum representations).
@@ II.1 A set of general initial states
The set of eigenstates of an infinite square well defined in $x\in[-{\sigma}/2,{\sigma}/2]$ is complete for functions whose primitive functions are piecewise continuous in $(-{\sigma}/2,{\sigma}/2)$ and have square-integrable first derivatives in their continuous regions[discrete].
@@
${\Psi}_{j}(x)=\begin{cases}\sqrt{\dfrac{2}{\sigma}}\sin\left(\dfrac{j\pi}{\sigma}(x-\tau)\right),&\tau<x<\tau+\sigma\\ 0,&\text{otherwise}\end{cases}.$ (1)
@@ II.2 Orthonormalization of free states
$I(\kappa-\kappa^{\prime})=\begin{cases}\text{finite large number},&\text{if }\kappa\neq\kappa^{\prime}\\ \infty,&\text{if }\kappa=\kappa^{\prime}\end{cases}.$ (2)
Where $x_{1},x_{2}$ are arbitrary finite coordinates and $\kappa(\varepsilon),\kappa^{\prime}(\varepsilon)$ are spectral function($\varepsilon$ is the continuous energy eigenvalue,proceed to the next subsection). Here, for ”orthonormalization”, instead of $\delta(\kappa-\kappa^{\prime})$, we use the symbol $I(\kappa-\kappa^{\prime})$. this is because the former is already employed in the context of completeness. For Eq. (2), one can understand it as follows: the average probability amplitude of the free state on both sides for ”normalization” is $\sqrt{1/(2\pi)}$, while the average probability amplitude of the free state restricted to one side only for ”normalization” is $\sqrt{1/\pi}$. Such Orthonormalization is derived from the ratio of infinities (or equivalently, Fourier transforms and sine-cosine transforms [Cos-sin]). Thus, the actual meaning of ”=” here is equivalent infinity symbol $\sim$”.
@@
Accounting for spatial symmetry and to avoid the effects of the order of limits, we adopt the Cauchy principal value integral [Cauchy].
@@ II.3 The spectral function
In general, $\kappa$ is independent of the potential at finite distances (Eqs. (5), (6)). For measurement normalization and the expansion of the initial state, $\kappa$ serves as the original integration variable (Eqs. (36), (22)). In different systems, $\kappa$ may take different forms, with the common ones being the wave number (Eqs. (6)) and the Bloch wave number [Bloch].
@@ III.1 Infinite potential on two sides
In a potential of this form, the energy eigenstates are bound states, and their completeness has been resolved [Hilbert]. For the completeness of solutions to the Sturm - Liouville equation, the sufficient condition the potential must satisfy is that there are finitely many discontinuities of finite magnitude in the potential. For continuous states, one can give an example that does not require this condition. The Dirac comb (with infinitely many discontinuities of infinite magnitude, Subsec. D) nevertheless has a set of complete energy eigenstates.
@@ III.2 Equal constant potential on two sides
In such a potential, the spectrum is continuous or mixed. The completeness of sets containing only pure continuous spectrum has been rigorously established via Fourier transforms (Dirichlet’s theorem [Dirichlet]). The universal SFs(let $m=\hbar=1$):
@@
There exist many examples of such SFs, and some of them have been proven. Examples proven for ${\kappa}(\varepsilon)=\sqrt{2\varepsilon}$ include reflectionless potentials [Reflectionless], $\delta$ potentials [delta], and Coulomb potentials [Coulomb], among others.
@@ III.2.1 The set of eigenstates
${\psi}_{n}^{1}(x)=A\begin{cases}e^{kx},&x<0\\ {(k/{k_{1}})\sin\left(k_{1}x\right)}+\cos\left(k_{1}x\right),&0<x<1\\ Be^{k_{2}(x-1)}+Ce^{-k_{2}(x-1)},&1<x<2\\ De^{-k(x-2)},&x>2\end{cases}.$ (9)
@@
${\psi}_{n}^{2}(x)=A^{\prime}\begin{cases}e^{kx},&x<0\\ {(k/{k_{3}})\sin\left(k_{3}x\right)}+\cos\left(k_{3}x\right),&0<x<1\\ B^{\prime}\cos\left(k_{4}(x-1)\right)+\\ C^{\prime}\sin\left(k_{4}(x-1)\right),&1<x<2\\ D^{\prime}e^{-k(x-2)},&x>2\end{cases}.$ (11)
@@
${\psi}_{\kappa}^{3}(x)=\begin{cases}\mathcal{A}e^{i\kappa x}+\mathcal{B}e^{-i\kappa x},&x<0\\ \mathcal{C}e^{ik_{5}x}+\mathcal{D}e^{-ik_{5}x},&0<x<1\\ \mathcal{E}e^{ik_{6}x}+\mathcal{F}e^{-ik_{6}x},&1<x<2\\ \mathcal{G}e^{i\kappa x},&x>2\end{cases}.$ (12)
@@
${\psi}_{\kappa}^{4}(x)=\begin{cases}\mathcal{G^{\prime}}e^{-i\kappa x},&x<0\\ \mathcal{E^{\prime}}e^{ik_{5}x}+\mathcal{F^{\prime}}e^{-ik_{5}x},&0<x<1\\ \mathcal{C^{\prime}}e^{ik_{6}x}+\mathcal{D^{\prime}}e^{-ik_{6}x},&1<x<2\\ \mathcal{A^{\prime}}e^{i\kappa x}+\mathcal{B^{\prime}}e^{-i\kappa x},&x>2\end{cases}.$ (15)
@@ III.3 Identical finite periodic potential on two-sides
We do not consider non-periodic oscillatory potentials at infinity, as such potentials are not well-suited to physical requirements, or equivalently, cannot be detected experimentally. Examples of periodic potentials include Bloch electrons [Bloch], Mathieu equations [Mathieu], among others. This section discusses two cases: the completeness of eigenstates in a cosine potential and the incompleteness of eigenstates in a discontinuous periodic potential.
@@ III.3.1 The set of eigenstates in a cosine potential
where $\text{MC}[2\varepsilon,1,x]$ and $\text{MS}[2\varepsilon,1,x]$ denote the even and odd Mathieu functions in Wolfram Mathematica, respectively. To uniquely determine the energy eigenstates, we may adopt the Bloch method [Bloch1] (or the Floquet method [Floquet]). By virtue of the potential’s periodicity $\pi$, the system possesses discrete translational symmetry, whence we obtain:
@@
where MS’ denotes the partial derivative of MS with respect to $x$. From Eq. 27, we note that $1<\kappa\leq 1$ or $1\leq\kappa<1$. With $\kappa$ restricted to this range, $(\kappa,\varepsilon)$ form a complete set of quantum numbers (with the eigenstates being orthonormal), or equivalently, this corresponds to the use of the reduced zone scheme [ReduceZone]. $\kappa(\varepsilon)$ is the inverse of the band function, whose band structure is illustrated in Fig. 2a-c.
@@ III.3.2 The completeness of the set of eigenstates in a cosine potential
For convenience, it is agreed that the integral sign $\int\limits_{0}^{\infty}$” denotes the integral excluding discontinuous parts within the integral range, and this convention shall apply to all subsequent contexts. From Figs. 2a–c, we observe that $\frac{\,d\kappa}{\,d\varepsilon}$ is infinite at the intersection points with the horizontal axis. This singularity is referred to as a Van Hove singularity [VanHove]. As inferred from Eqs. 31 and 32, the integral exists but is difficult to compute accurately. As shown in Figs. 2e–g, for $j=1$, ${f}_{j}(x)$ deviates from $\Psi_{j}(x)$ by up to 0.012. However, for $j=2$ and $3$, the completeness exhibits excellent behavior. Furthermore, as we will see in the next subsection (the Dirac comb), the spectral function and completeness perform well in periodic potentials. Thus, the deviation for $j=1$ is an algorithmic issue and does not affect our conclusions.
@@ III.3.3 The incompleteness of eigenstates in periodic square potential
$V\left(x\right)=\begin{cases}{V}_{0},&-b<x<0\\ {V}_{1},&0<x<1\end{cases}.$
@@
Solutions to the eigenstates of periodic square potential (Kronig-Penney model) can be found in some quantum mechanics textbooks [periodicsquare]; we do not elaborate on them here. To demonstrate the incompleteness of its eigenstates, it suffices to show that the measurement probability does not normalize to unity—since the normalization of measurement probability is a necessary condition for the completeness of eigenstates [Landau]. The measurement probability amplitudes of the general initial state (Eq.1) in the above four cases are:
@@ III.3.4 A physically valid initial state
$\mathcal{V}\left(x\right)=\begin{cases}+\infty,&x<\tau\\ {V}_{0},&\tau<x<0\\ {V}_{1},&0<x<\tau+\sigma\\ +\infty,&x>\tau+\sigma\end{cases}.$ (37)
@@
$\phi(x)=A\begin{cases}0,&x<\tau\\ \sin(\xi(x-\tau)),&\tau<x<0\\ \frac{\sin(\xi\tau)}{\sin(\eta(\tau+\sigma))}\sin(\eta(\tau+\sigma-x)),&0<x<\tau+\sigma\\ 0,&x>\tau+\sigma\end{cases}.$ (38)
@@
Using the eigenstate set from the previous subsubsection to expand the initial state in Eq. 38, we find that the total measurement probability is $P=0.417$. This indicates the incompleteness of eigenstates in periodic potentials with finite discontinuities, or equivalently, in infinite sets of finite-discontinuity potentials—consistent with the S-L theorem for purely discrete spectra [discrete]. In contrast, as will be shown in the next subsubsection, the eigenstates in the Dirac comb potential exhibit completeness.
@@ III.4.1 The set of eigenstates in the Dirac comb potential
$\psi(x)=\mathcal{A}\begin{cases}\mathcal{B}e^{-ikx}+e^{ikx},&-a/2<x<0\\ e^{ia\kappa}\left(\mathcal{B}e^{-ik(x-a)}+e^{ik(x-a)}\right),&0<x<a/2\end{cases}.$ (40)
@@ III.4.2 The completeness of the set of eigenstates in the Dirac comb potential
Here,$a$ is the period. Many properties of solutions to the Mathieu equation are shared by solutions to the Hill equation[Hill].
@@ IV.1 Unequal constant potential on two sides
$\mathcal{V}(x)=\begin{cases}-V_{0}/2,&x<0\\ V_{0}/2,&x>0\end{cases}.$ (45)
@@
$V\left(x\right)=\begin{cases}0,&x<0\\ V_{0},&x>0\end{cases}.$ (46)

### v2

@@ Abstract
The Schrödinger equation (SE) is the fundamental dynamical equation describing quantum systems, and its linearity directly gives rise to the superposition principle. The completeness of a set of orthonormal eigenstates precisely characterizes this superposition principle. Completeness requires that any state of the system can be expressed as a linear combination of these eigenstates; conversely, any linear combination of the eigenstates also constitutes a possible state of the system [22]. This property has been intuitively verified in numerous Young’s double-slit experiments [8] and enables the prediction of quantum jump probabilities [19].Born’s probability interpretation [4] further endows eigenstates with physical significance: the squared modulus of the coefficient of each eigenstate in a superposition corresponds to the probability of the particle being in that state. The physical meaning of completeness is interpreted as a representation transformation [23].
@@
The time-independent SE is a type of (generalized) Sturm-Liouville (S-L) equation. Mathematical research on the completeness of solutions to S-L equations dates back to the Hilbert-Schmidt theory in 1907 [12]. In 1910, Weyl extended S-L equations from finite to infinite intervals. Courant and Hilbert proved the completeness of the discrete spectrum (bound states) in 1924 [5]. Titchmarsh systematically developed the eigenfunction expansion theory for singular S-L operators in 1946, establishing the completeness of the continuous spectrum (free states) and mixed spectrum (coexisting bound and free states) under certain conditions [28]. Kodaira comprehensively summarized previous results in 1949 and developed simplified proof methods for completeness [15][9]. Since then, progress in completeness research has been slow—no general proof for the completeness of S-L equations has been achieved to date. Additionally, mathematical proofs of completeness are complex and lack clear physical interpretation.
@@
For quantum mechanics, however, it suffices to prove the completeness of SE solutions. Furthermore, the potential function can be restricted to forms encountered in physical scenarios, eliminating the need for generalized discussions. We first classify potential functions into two categories: symmetric and asymmetric potentials (Sec. III and IV). Following Titchmarsh’s classification method [29], we further divide these two categories into 8 types based on four possible asymptotic behaviors of the potential at infinity (infinite, constant, finite oscillatory, and infinite oscillatory). Among these 8 types, only one has an existing general theoretical proof (Subsection IIIA). For the remaining cases, we provide theoretical proofs for some and conduct completeness tests via numerical simulations for others. In the course of these proofs, the spectral function retains physical meaning (momentum spectrum), and the form of completeness aligns with the interpretation of representation transformation (coordinate and momentum representations).
@@ II.1 A set of general initial states
The set of eigenstates of an infinite square well defined in $x\in[-{\sigma}/2,{\sigma}/2]$ is complete for functions whose primitive functions are piecewise continuous in $(-{\sigma}/2,{\sigma}/2)$ and have square-integrable first derivatives in their continuous regions[5].
@@
[…]
@@ II.2 Orthonormalization of free states
$I(\kappa-\kappa^{\prime})=\cases{}{finitelargenumber},&\text{if }\kappa\neq\kappa^{\prime}\\ \infty,&\text{if }\kappa=\kappa^{\prime}.$ (2)
Where $x_{1},x_{2}$ are arbitrary finite coordinates and $\kappa(\varepsilon),\kappa^{\prime}(\varepsilon)$ are spectral function($\varepsilon$ is the continuous energy eigenvalue,proceed to the next subsection). Here, for ”orthonormalization”, instead of $\delta(\kappa-\kappa^{\prime})$, we use the symbol $I(\kappa-\kappa^{\prime})$. this is because the former is already employed in the context of completeness. For Eq. (2), one can understand it as follows: the average probability amplitude of the free state on both sides for ”normalization” is $\sqrt{1/(2\pi)}$, while the average probability amplitude of the free state restricted to one side only for ”normalization” is $\sqrt{1/\pi}$. Such Orthonormalization is derived from the ratio of infinities (or equivalently, Fourier transforms and sine-cosine transforms [26]). Thus, the actual meaning of ”=” here is equivalent infinity symbol $\sim$”.
@@
Accounting for spatial symmetry and to avoid the effects of the order of limits, we adopt the Cauchy principal value integral [30].
@@ II.3 The spectral function
In general, $\kappa$ is independent of the potential at finite distances (Eqs. (5), (6)). For measurement normalization and the expansion of the initial state, $\kappa$ serves as the original integration variable (Eqs. (36), (22)). In different systems, $\kappa$ may take different forms, with the common ones being the wave number (Eq. (6)) and the Bloch wave number [13].
@@ III.1 Infinite potential on two sides
In a potential of this form, the energy eigenstates are bound states, and their completeness has been resolved [12]. For the completeness of solutions to the Sturm - Liouville equation, the sufficient condition the potential must satisfy is that there are finitely many discontinuities of finite magnitude in the potential. For continuous states, one can give an example that does not require this condition. The Dirac comb (with infinitely many discontinuities of infinite magnitude, Subsec. D) nevertheless has a set of complete energy eigenstates.
@@ III.2 Equal constant potential on two sides
In such a potential, the spectrum is continuous or mixed. The completeness of sets containing only pure continuous spectrum has been rigorously established via Fourier transforms (Dirichlet’s theorem [25]). The universal SFs(let $m=\hbar=1$):
@@
There exist many examples of such SFs, and some of them have been proven. Examples proven for ${\kappa}(\varepsilon)=\sqrt{2\varepsilon}$ include reflectionless potentials [7], $\delta$ potentials [24], and Coulomb potentials [20], among others.
@@ III.2.1 The set of eigenstates
${\psi}_{n}^{1}(x)=A\cases{e}^{kx},&x<0\\ {(k/{k_{1}})\sin\left(k_{1}x\right)}+\cos\left(k_{1}x\right),&0<x<1\\ Be^{k_{2}(x-1)}+Ce^{-k_{2}(x-1)},&1<x<2\\ De^{-k(x-2)},&x>2.$ (9)
@@
${\psi}_{n}^{2}(x)=A^{\prime}\cases{e}^{kx},&x<0\\ {(k/{k_{3}})\sin\left(k_{3}x\right)}+\cos\left(k_{3}x\right),&0<x<1\\ B^{\prime}\cos\left(k_{4}(x-1)\right)+\\ C^{\prime}\sin\left(k_{4}(x-1)\right),&1<x<2\\ D^{\prime}e^{-k(x-2)},&x>2.$ (11)
@@
${\psi}_{\kappa}^{3}(x)=\cases{}{A}e^{i\kappa x}+\mathcal{B}e^{-i\kappa x},&x<0\\ \par\mathcal{C}e^{ik_{5}x}+\mathcal{D}e^{-ik_{5}x},&0<x<1\\ \par\mathcal{E}e^{ik_{6}x}+\mathcal{F}e^{-ik_{6}x},&1<x<2\\ \par\mathcal{G}e^{i\kappa x},&x>2.$ (12)
@@
${\psi}_{\kappa}^{4}(x)=\cases{}{G^{\prime}}e^{-i\kappa x},&x<0\\ \par\mathcal{E^{\prime}}e^{ik_{5}x}+\mathcal{F^{\prime}}e^{-ik_{5}x},&0<x<1\\ \par\mathcal{C^{\prime}}e^{ik_{6}x}+\mathcal{D^{\prime}}e^{-ik_{6}x},&1<x<2\\ \par\mathcal{A^{\prime}}e^{i\kappa x}+\mathcal{B^{\prime}}e^{-i\kappa x},&x>2.$ (15)
@@ III.3 Identical finite periodic potential on two-sides
We do not consider non-periodic oscillatory potentials at infinity, as such potentials are not well-suited to physical requirements, or equivalently, cannot be detected experimentally. Examples of periodic potentials include Bloch electrons [13], Mathieu equations [10], among others. This section discusses two cases: the completeness of eigenstates in a cosine potential and the incompleteness of eigenstates in a discontinuous periodic potential.
@@ III.3.1 The set of eigenstates in a cosine potential
where $\text{MC}[2\varepsilon,1,x]$ and $\text{MS}[2\varepsilon,1,x]$ denote the even and odd Mathieu functions in Wolfram Mathematica, respectively. To uniquely determine the energy eigenstates, we may adopt the Bloch method [27] (or the Floquet method [6]). By virtue of the potential’s periodicity $\pi$, the system possesses discrete translational symmetry, whence we obtain:
@@
where MS’ denotes the partial derivative of MS with respect to $x$. From Eq. 27, we note that $1<\kappa\leq 1$ or $1\leq\kappa<1$. With $\kappa$ restricted to this range, $(\kappa,\varepsilon)$ form a complete set of quantum numbers (with the eigenstates being orthonormal), or equivalently, this corresponds to the use of the reduced zone scheme [14]. $\kappa(\varepsilon)$ is the inverse of the band function, whose band structure is illustrated in Fig. 2a-c.
@@ III.3.2 The completeness of the set of eigenstates in a cosine potential
For convenience, it is agreed that the integral sign $\int\limits_{0}^{\infty}$” denotes the integral excluding discontinuous parts within the integral range, and this convention shall apply to all subsequent contexts. From Figs. 2a–c, we observe that $\frac{\,d\kappa}{\,d\varepsilon}$ is infinite at the intersection points with the horizontal axis. This singularity is referred to as a Van Hove singularity [17]. As inferred from Eqs. 31 and 32, the integral exists but is difficult to compute accurately. As shown in Figs. 2e–g, for $j=1$, ${f}_{j}(x)$ deviates from $\Psi_{j}(x)$ by up to 0.012. However, for $j=2$ and $3$, the completeness exhibits excellent behavior. Furthermore, as we will see in the next subsection (the Dirac comb), the spectral function and completeness perform well in periodic potentials. Thus, the deviation for $j=1$ is an algorithmic issue and does not affect our conclusions.
@@ III.3.3 The incompleteness of eigenstates in periodic square potential
$V\left(x\right)=\cases{V}_{0},&-b<x<0\\ {V}_{1},&0<x<1.$
@@
Solutions to the eigenstates of periodic square potential (Kronig-Penney model) can be found in some quantum mechanics textbooks [11]; we do not elaborate on them here. To demonstrate the incompleteness of its eigenstates, it suffices to show that the measurement probability does not normalize to unity—since the normalization of measurement probability is a necessary condition for the completeness of eigenstates [16]. The measurement probability amplitudes of the general initial state (Eq.1) in the above four cases are:
@@ III.3.4 A physically valid initial state
$\mathcal{V}\left(x\right)=\cases{+}\infty,&x<\tau\\ {V}_{0},&\tau<x<0\\ {V}_{1},&0<x<\tau+\sigma\\ +\infty,&x>\tau+\sigma.$ (37)
@@
$\phi(x)=A\cases{0},&x<\tau\\ \sin(\xi(x-\tau)),&\tau<x<0\\ \frac{\sin(\xi\tau)}{\sin(\eta(\tau+\sigma))}\sin(\eta(\tau+\sigma-x)),&0<x<\tau+\sigma\\ 0,&x>\tau+\sigma.$ (38)
@@
Using the eigenstate set from the previous subsubsection to expand the initial state in Eq. 38, we find that the total measurement probability is $P=0.417$. This indicates the incompleteness of eigenstates in periodic potentials with finite discontinuities, or equivalently, in infinite sets of finite-discontinuity potentials—consistent with the S-L theorem for purely discrete spectra [5]. In contrast, as will be shown in the next subsubsection, the eigenstates in the Dirac comb potential exhibit completeness.
@@ III.4.1 The set of eigenstates in the Dirac comb potential
$\psi(x)=\mathcal{A}\cases{}{B}e^{-ikx}+e^{ikx},&-a/2<x<0\\ e^{ia\kappa}\left(\mathcal{B}e^{-ik(x-a)}+e^{ik(x-a)}\right),&0<x<a/2.$ (40)
@@ III.4.2 The completeness of the set of eigenstates in the Dirac comb potential
Here,$a$ is the period. Many properties of solutions to the Mathieu equation are shared by solutions to the Hill equation[18].
@@ IV.1 Unequal constant potential on two sides
$\mathcal{V}(x)=\cases{-}V_{0}/2,&x<0\\ V_{0}/2,&x>0.$ (45)
@@
$V\left(x\right)=\cases{0},&x<0\\ V_{0},&x>0.$ (46)